OpenStudy (anonymous):

find the slope-intercept form of the line that passes through the point (-6,4) and is parallel to the line 2x-2y=6

OpenStudy (anonymous):

if two lines are parallel, then they have the same slope

OpenStudy (anonymous):

yes i dont quite remember how to do this kind aof problem

OpenStudy (anonymous):

Let's rewrite the equation for the given line in slope intercept form as it will make things easier. 2x - 2y = 6 2x - 6 = 2y y = x - 3 so, we know from that equation that the slope of the line we are after is 1 as: y = (1 * x) - 3 and we are given that our new line shares this slope ( is parallel to it ) New Line: y = 1x +b Now, we also know that the new line passes a given point (-6, 4). so, we now can sub in those values and solve for "B"